A courier service company has found that their delivery time of parcels are 50 minutes and variance of 25 minutes squared
What is the probability that a randomly selected parcel will take 60 minutes to deliver?
1
Expert's answer
2020-06-11T20:46:28-0400

Mean delivery time = 50 minutes


Variance = 25 minutes




probability that a randomly selected parcel will take 60 minutes to deliver



Variance = 25 minutes


Therefore, Standard deviation "= \sqrt{25} =5"


"Z score = \frac{( Given \, value - Mean)}{Standard \,deviation}"


Z score for 60 "=\frac{ (60 - 50)} {5} = 2"


"Z(2) =" 0.97725